OpenStudy (anonymous):

[6.03] What is the value of the x variable in the solution to the following system of equations? 4x + 3y = 10 4x + 5y = 14 Select one: a. 2 b. 1 c. -1 d. -2

OpenStudy (jadeishere):

Substitute the values that are given into the equation... 4(1) + 3y = 10 4(1) + 5y = 14 4 + 3y = 10 4 + 5y = 14 3y = 6 5y = 10 x = 1 I'm pretty sure

OpenStudy (anonymous):

Thank you so much! @Jadeishere

OpenStudy (jadeishere):

Yep :)